Which Professor Is Better Off? Suppose the starting salary for a new assistant economics professor was $20,000 in 1976 and $92,000 in 2014. The value of the CPI for 2014 was 236.7, compared to 56.9 in 1976. A newly-hired professor earned more in rea tmwith a real salary of S(Enter your response rounded to the nearest integer.)

(The First Drop down arrow box gives you 1976 and 2014 as an option)
Answer #1

Correct Answer:

2014

$22115.76 or $22116

Working note:

Real value of $20000 in 2014 = 20000*(CPI in 2014/CPI in 1976) =  20000*(236.7/56.9) = $83198.59

Since, professor in 2014 is getting 92000 that is higher than $83198.59. So, professor is better off in 2014.

Real value of $92000 in 2014 = 92000*(56.9/236.7) = $22115.76 or $22116